# DexCatch: Learning to Catch Arbitrary Objects with Dexterous Hands

### v1

@@ III-B Task Desciption
The 3D point cloud feature of an object represents its shape, created by densely gathered points in 3D space. We adopt point clouds to approximate object features. Our approach involves feature extraction via a point clouds processing network utilizing PointClipV2 [24]. This algorithm facilitates zero-shot classification segmentation to identify features. To reduce computational complexity, the extracted feature map degrades to two dimensions using Principal Component Analysis (PCA). Our experimentation and visual analysis have substantiated that these 2-dimensional features adequately capture the distinct attributes of various objects, as demonstrated in Fig.2. It can be seen that the distribution of various objects extends along three distinct directions. The phenomenon allows for the classification of objects into three categories. The first category comprises elongated objects, exemplified by items like bottles. The second category encompasses flat objects, such as keyboards. Lastly, regular objects, which do not fit into the previous categories, such as cups, are classified under the third category. Furthermore, we add variations in the mass and inertia of objects during training, leading to the generalization ability about the different physical properties of objects.
TABLE I: Reward Desgin
@@
State, Action and Reward
@@
1) State: The state space consists of 424 elements. Among these, the 422-dimensional state incorporates crucial information such as hand position, hand orientation, angular and linear velocities of the hands, position and speed of each hand’s degree of freedom, force-torque sensor readings of the hands, target object’s position and orientation, as well as the current object’s position and orientation. The above observation corresponds to the setting in [22]. Additionally, a two-dimensional state features the input derived from the point cloud data of the object. It serves as the perceptual part for the agent regarding the object. 2) Action: The action space comprises 52 dimensions, encompassing 20 pairs of agonist-antagonist tendons for both hands, along with three dimensions each for the required force and torque of each hand. 3) Reward: The reward is composed of five elements: 1) $r_{p}$ quantifies the distance between the position of the goal object $p_{g}$ and the current object $p_{c}$. 2) $r_{o}$ represents the angular discrepancy between the target object’s orientation and the current object’s orientation, utilizing quaternions for accurate angular measurement. 3) $r_{a}$ evaluates the smoothness of joint movement through the current actions $a_{c}$. 4) A constant reward is given for successfully completing tasks $r_{succ}$. 5) A penalty is imposed for task failure $r_{fall}$. Specifically, the smoothness of joint movement is controlled by negatively rewarding the sum of squares of action quantities as per the strategy. The overall reward is expressed $r_{total}$ in Table I, where target position distance is calculated using Euler distance, angular distance is determined using quaternions, joint smoothness is penalized using action quantities’ sum of squares, and success and failure rewards are assigned as positive and negative constants, as outlined in Table I.
IV Stability-Constrained Reinforcement Learning Algorithm
To address the task outlined in the above section, we propose the Stability-Constrained Reinforcement Learning (SCRL) algorithm grounded in the Proximal Policy Optimization (PPO) [26] framework. Our approach integrates considerations for both system stability during task completion and efficient learning. Furthermore, we incorporate parameter space normalization to speed up the learning.
The algorithm’s architecture is illustrated in Fig. 3. Here’s a breakdown of its components:1) Point Cloud Feature Extraction: The initial step involves extracting point cloud features using the PointClipV2 [24]. Subsequently, Principal Component Analysis (PCA) is employed to reduce dimensionality, yielding a 2-dimensional feature output. This output serves as the perception state for the shadow hands system; 2) Observation Input: The agent takes both the current observation of objects and the shadow hands and the compressed features from point clouds as input; 3) Actor-Critic Framework: The actor and critic are represented as two neural networks, which is the same as PPO algorithm. The actor takes the current state $s_{t}$ as input and executes the action $a_{t}$. The system interacts with the environment to obtain the reward $r_{t}$ and the next candidate state $s^{\prime}_{t+1}$. The critic evaluates the candidate state $s^{\prime}_{t+1}$ to estimate its value $v^{\prime}$. Meanwhile, a Lyapunov function takes as input $s_{t}$, yielding the Lyapunov value $v_{L}$ for the current state; 4) Replay Buffer and Parameter Update: Relevant data from the current state is stored in the replay buffer. The total advantage $A_{all}$ is calculated based on $r_{t}$, $v$, $v^{\prime}$, $v_{L}$ and $v^{\prime}_{L}$, which plays a role in updating the parameters of the actor. The specific formulation of $A_{all}$ is denoted in Eq. (10). The synergy of these steps constitutes the SCRL algorithm, achieving both stability and efficient exploration during task execution.
IV-A Proximal Policy Optimization(PPO)
The PPO algorithm is an on-policy reinforcement learning approach. It excels in robotics tasks featuring high-dimensional action spaces. The algorithm contains an actor and a critic network, where $\theta$ and $\phi$ denote the parameters of the actor network $\pi_{\theta}(a_{t}|s_{t})$ and the critic network $V_{\phi}(s_{t})$, respectively. For the policy function $\pi_{\theta}(a_{t}|s_{t})$, the output corresponds to the mean and std of Gaussian distribution. On the other hand, the value function $V_{\phi}(s_{t})$ yields the current state’s value. Building upon the policy gradient method, the parameter updating formula for the actor network takes the following form:
@@ IV-A Proximal Policy Optimization(PPO)
$L(\cdotp)=\min(\tilde{A}(\boldsymbol{s},\boldsymbol{a}),\emph{clip}(1,1-\epsilon,1+\epsilon)\tilde{A}(\boldsymbol{s},\boldsymbol{a}))$ (3)
$\tilde{A}(\boldsymbol{s},\boldsymbol{a})=\frac{\pi_{\theta}(\boldsymbol{a}|\boldsymbol{s})}{\pi_{\theta_{k}}(\boldsymbol{a}|\boldsymbol{s})}A^{\pi_{\theta_{k}}(\boldsymbol{s},\boldsymbol{a})}$ (4)
where ${A^{\pi_{\theta_{k}}}(\boldsymbol{s},\boldsymbol{a})=G^{\pi_{\theta_{k}}}_{t}-V^{\pi_{\theta_{k}}}_{\phi}(s_{t})}$, is the advantage function, represents the tolerant range of deviation between the original and existing strategy. For parameters $\phi$, the update is expressed as:
$\begin{array}[]{l}\phi_{k+1}=\underset{\phi}{\arg max}\underset{\boldsymbol{s,a}\sim\pi_{\theta_{k}}}{E}[G^{\pi_{\theta_{k}}}_{t}-V^{\pi_{\theta_{k}}}_{\phi}(s_{t})]\\ \end{array}$ (5)
However, because the calculation of $G^{\pi_{\theta_{k}}}_{t}$ value needs the whole data of a finished episode, we approximate $G^{\pi_{\theta_{k}}}_{t}$ to $r_{t}+V^{\pi_{\theta_{k}}}_{\phi}(s_{t+1})$ through TD learning method.
IV-B Lyapunov Stability
The Lyapunov function serves as a potent tool for assessing system stability. In the context of RL, the Markov Decision Process (MDP) represents a discrete-time dynamical system corresponding to continuous states, and the Lyapunov stability condition becomes a constraint within the objective function. This integration ensures both stability and optimality of the system. Indeed, it’s important to note that in our framework, the points of optimality and stability do not always coincide. As depicted in Fig. 4, the middle-left figure illustrates a gesture that is optimal for the task but lacks stability, whereas the right figure showcases a gesture that combines both stability and optimality.
One approach, known as the self-learning Lyapunov stability method [27], is utilized in our paper. First, we introduce a Lyapunov stability critic network, which represents the candidate Lyapunov function $V_{L}$ using neural networks. Subsequently, the empirical Lyapunov risk $\mathcal{R}(\Lambda)$ [27] is defined as per Eq. (6). The parameter update is then executed by minimizing the loss function associated with the Lyapunov risk $\mathcal{R}(\Lambda)$, as part of achieving the desired system stability.
$\mathcal{R}(\Lambda)$ $=\frac{1}{N}\sum^{N}_{i=1}(\max(-V_{L}(s_{i}),0)$ (6)
$+\max(0,L_{f}V_{L}))+V_{L}^{2}(0))$
where $s_{1},\cdots,s_{i},\cdots,s_{n}$ is the state of sampling. The empirical Lyapunov risk is non-negative, we approximate the Lie derivative $L_{f}V_{L}$ by finite difference of the sampling trajectory of the system because of the complexity of modeling the dynamics of this system:
$\begin{array}[]{cc}L_{f}V_{L}(s)=\lim_{\Delta t\to 0}\frac{1}{\Delta t}(V_{L}(s^{\prime})-V_{L}(s))\end{array}$ (7)
where $s$ and $s^{\prime}$ are two consecutive states, and $\Delta t$ is the time difference between them. Since the Lyapunov function is considered to be another value function, Lie derivative term $L_{f,\Delta t}V_{L}(s)$ is the only term in the Lyapunov risk affected by the policy. Therefore, an additional goal of ensuring $L_{f,\Delta t}V_{L}(s)<0$ is added to the policy optimization, to encourage the policy’s update with the requirement of stabilization. Hereafter, we design the part of the Lyapunov stability advantage function based on the clipped Lie derivative term:
$A_{L}(s_{t},a_{t})=\min(0,-L_{f_{\pi_{\theta}},\Delta t}V_{L}(s_{t}))$ (8)
Figure 4: The solution with stability and the optimal solution are not consistent in the task. For example, the optimal solution on the left, the reward and advantage remain the highest but the gesture is not stable. In the stable solution on the right, the hand totally hold the object.
Figure 5: Video snapshots of throwing and catching tasks performed by our method. There exist three tasks with different postures of hands, Overarm Catch, Abreast Catch and Underarm Catch. The most challenging task is Overarm Catch because the hands are oriented vertically. Notably, our method can catch diverse objects in three tasks.
IV-C Intrinsic Advantage
Based on the advantage function of the PPO algorithm, in order to improve the ability of the system efficiently, we design the intrinsic advantage function with reference to the intrinsic reward mechanism [28]. For a Markov decision process trajectory, different values $V_{\phi}(s_{t})$ will affect the success of the final result of this trajectory. To provide a specific example, let’s consider the fact that the next states $s^{\prime}_{t+1}$ exhibit varying conditions, leading to different values $V_{\phi}(s_{t+1})$. It’s worth noting that in a successful process, the value $V_{\phi}(s_{t+1})$ tends to be higher. Consequently, within a Markov decision process, we can further refine the advantage function $A(s,a)$ using empirical measures. If an action is effective, it should result in a higher value, as represented by $V_{\phi}(s_{t+1})-V_{\phi}(s_{t})>0$. This signifies that action $a_{t}$ transitions to a state with a higher success rate. Conversely, an unfavorable action can lead to a decrease in value, reflected as $V_{\phi}(s_{t+1})-V_{\phi}(s_{t})<0$. In order to reduce failure in the early stage of the throwing process, we consider the intrinsic advantage $A_{I}(s_{t},s_{t+1})$ as part of the total advantage function to accelerate the RL training, shown as follows:
IV-D Network Structure
The network architecture for the Lyapunov function is depicted on the right side of Fig. 3. This structure comprises three linear layers, with a Batch normalization layer and an ELU activation function appended to the output of each layer. Following these, there is a penultimate normalization layer [29], succeeded by a linear layer to regulate the output dimension. The network structure for the policy and value function are similar to that of the Lyapunov value network.
To sum up, the total advantage $A_{all}$ consists of 3 parts: standard advantage estimation based on PPO, intrinsic advantage estimation, and stability advantage estimation. The formula for the total advantage is expressed as follows:
$\begin{array}[]{cc}A_{all}(s_{t},a_{t},s_{t+1})\\ =A^{\pi_{\theta}}(s_{t},a_{t})+\beta_{1}A_{I}(s_{t},s_{t+1})+\beta_{2}A_{L}(s_{t},a_{t})\\ \end{array}$ (10)
where $A^{\pi_{\theta}}(s_{t},a_{t})$ refers to the time differential advantage function of the PPO algorithm, $A_{I}(s_{t},s_{t+1})$ represents the intrinsic advantage function, and $\beta_{1},\beta_{2}\in[0,1]$ are hyperparameters, $A_{L}(s_{t},a_{t})$ represents the Lyapunov stability advantage function.
V EXPERIMENTS
To test the algorithm’s ability to generalize, we trained and tested it with objects from the YCB__\__Pybullet dataset [30]. The results show that our algorithm outperforms baseline methods. Four essential components of our algorithm significantly influence its performance. Additionally, our method showcases the ability to generalize different gestures across three skillful throwing and catching tasks, effectively allowing for the successful manipulation of various objects. Importantly, our approach extends its capabilities to successfully catch previously unseen objects.
Figure 6: The average reward curve for training multiple objects for three different tasks. The first row is the comparison experiment reward graph. The second row is an ablation experiment reward graph. (Curves smoothed report the mean 5 times, and shadow areas show the variance.)
TABLE II: Success rate of Task
@@ V EXPERIMENTS
Train Test
Apple Banana Peach Pen Strawberry Pie Poker Sugar Mug META Stapler Scissor Washer Bowl META
Overarm 42.55 53.82 37.20 68.30 28.56 52.71 41.77 48.41 32.20 45.05 23.44 28.40 76.28 82.22 52.58
Abreast 62.11 88.65 46.51 47.12 50.89 99.77 86.34 93.33 76.68 72.37 94.84 27.59 83.94 83.59 72.49
Underarm 71.51 38.80 73.28 31.81 44.62 69.22 42.62 26.34 76.26 52.71 71.33 31.80 57.67 53.24 53.51
V-A Comparison Experiments
In this section, we conducted a comparative analysis of our algorithm against mainstream reinforcement learning (RL) algorithms in dexterous hands-throwing and catching tasks. These algorithms include TRPO [31], SAC [32], DDPG [33], and PPO [26]. The results, as illustrated in Fig. 6, are quite telling. In the most challenging task, the Overarm Catch with vertical hands, our algorithm significantly outperforms the baseline methods, achieving success where others fail. For the Abreast Catch task, our approach demonstrates superior learning efficiency in acquiring throwing and catching skills. Furthermore, even in the relatively easier Underarm Catch task, our method exhibits slightly improved learning efficiency and performance compared to the baseline algorithms. These findings highlight the robustness and effectiveness of our algorithm, particularly in demanding scenarios. In the Underarm Catch task, as depicted in Fig. 5, the distance between the two hands is the shortest. Consequently, both the PPO and DDPG algorithms manage to solve this task, with DDPG even surpassing our algorithm in terms of average reward. However, it’s essential to note that during testing, DDPG and PPO produce less favorable results compared to our method. This discrepancy can be attributed to the fact that their strategies generate unstable grasping gestures, primarily due to the absence of stability requirements in their approaches. Our algorithm, with its focus on stability, excels in ensuring consistent and reliable grasping gestures, even in challenging scenarios. In summary, our algorithm exhibits superior learning efficiency and stability. Video snapshots illustrating the performance in three tasks can be seen in Fig. 5. It’s worth emphasizing that we trained the model using nine objects with diverse shapes and mass distributions. This showcases the algorithm’s robust generalization ability, as it can effectively adapt to various objects and perform tasks efficiently and reliably.
V-B Ablation Experiment
We evaluated the impact of four crucial components in our algorithm. To clarify, here are their definitions: $a.$ includes point cloud features from different objects as input. $b.$ represents the Lyapunov stability component. $c.$ stands for the intrinsic advantage component. $d.$ indicates parameter normalization. The effect of these components is shown in the reward curve in Figure 6. Our initial experiment involved the removal of component $a.$, which entails excluding the point cloud features of objects as input. The outcome demonstrated a substantial decrease in the average reward, particularly in the challenging Overarm Catch task. This result underscores the importance of providing the system with prior information about objects to facilitate faster learning in demanding tasks. In our next experiment, we took out components $a.$ and $b.$ from the algorithm. The reward curve showed a significant effect of the Lyapunov stability component on early learning and how quickly the algorithm converges. This part adds constraints to the learning process, ensuring that the dexterous hands focus on learning stable gestures for catching objects rather than just pursuing the highest rewards. This balance between stability and optimization is essential for the system’s performance. For training w.o. $a.$ $b.$ $c.$, it can be seen that the effect of the intrinsic advantage part is slight in Overarm Catch and Abreast Catch. Nevertheless, it offers an important advantage in performance in Underarm Catch. The experiment where we removed all components ($a.$, $b.$, $c.$, and $d.$) underscores the significance of parameter normalization. Notably, this experiment reveals that each element in our algorithm has a slight impact on the Underarm Catch task. This observation aligns with the understanding that the Underarm Catch task is relatively easier compared to the other tasks.
V-C Generalization Ability
For diverse objects: In this section, we delve into a detailed analysis of our method’s generalization capabilities. It demonstrates the ability to execute distinct catching gestures based on the positions of the shadow hands and the objects being caught. Through an analysis of the high-dimensional point cloud features, we categorized most objects into three primary classes. Remarkably, our method exhibits over three different catching gestures for each of these three object types, as depicted in Fig. 7. For objects falling into the regular, flat, and elongated categories, the model generalizes three distinct gestures in each of the three tasks. For instance, when dealing with regular objects like apples, it employs either a half-grip or pinching gesture. Slender objects like bananas are handled with a full-grip position. Flat objects, such as scissors and pokers, are directly held and secured between the fingers and palms. This versatility in catching gestures reflects the algorithm’s remarkable adaptability to different object types and positions.
For unseen objects: Moreover, we evaluate our method in the throwing and catching task using various objects, quantifying the success rate as a key metric. We run the testing task in 10k episodes and use the average reward threshold to determine whether it is successful. Since our algorithm pays more attention to the efficiency and stability of generalization ability learning, we no longer use the distance threshold of the final state from the target position as the success rate. The success rate of 13 objects is illustrated in the Table II. Among them, we use 4 unseen objects (test set) to verify the generalization ability of our method. Specifically, we adopt two methods to test objects in the training and test sets: i. Test the success rate of multiple objects in parallel. ii. Test the success rate of each object individually. We take the most challenging task, Overarm Catch as an example. The average success rate for the training set objects reaches $45.05\%$, and the highest of individual tests on each object was pen up to $68.30\%$. In addition, the average success rate of the test set object is close to $52.58\%$. Most importantly, our method obtains a great success rate, $28.40\%$, in testing scissors, which is a new object. Furthermore, we can observe that our method can provide a robust catching strategy because there is not a significant gap between the success rates of training and test objects. For Abreast Catch and Underarm Catch, our method achieves higher success rates in single and multiple objects tests. The results provide strong evidence of our algorithm’s robust generalization capability for catching diverse objects.
V-D Implementation detail
In our experiments, we conducted training and testing on a computing platform equipped with an RTX A6000 GPU and an AMD CPU. We utilized nine objects for training across three different task types and tested the algorithm with four objects. Our approach also leveraged meta-learning for multi-object parallel training. The objects used in our experiments were sourced from the YCB dataset [30]. Due to the limitations of the training platform, we employed 256 parallel environments per object for the Overarm Catch and Underarm Catch tasks. For the Abreast Catch task, each object was trained with 64 parallel environments. The policy updates occurred at a frequency of 0.125 times per environment step, and the length of each episode was set to 100 steps.
VI Conclusion
In this paper, we propose the Stability Constraint Reinforcement Learning (SCRL) algorithm to tackle the dexterous shadow-handed throwing and catching task across diverse objects. We evaluate its performance in Underarm Catch, Abreast Catch and Overarm Catch and find that the SCRL algorithm significantly outperforms baselines. Notably, our learned strategy exhibits remarkable zero-shot transfer capabilities when dealing with previously unseen objects. Even in challenging scenarios where the hand is oriented vertically, our algorithm excels in baselines by a large margin. The incorporation of Lyapunov stability constraints plays a pivotal role in enhancing the stability of the throwing and catching actions. The intrinsic advantage and the regularization of the parameter space mechanism accelerate early learning, leading to a higher performance. Our method autonomously generates different catching gestures for slender, flat, and regular objects. Additionally, it creates diverse catching behaviours when faced with different postures of a type of object. This research opens up promising avenues for the generalization of dynamic manipulation using dexterous hands, showcasing the potential of the RL-based approach to improve object transportation efficiency through throwing and catching.

### v2

@@ Appendix B Object-generalized design
The 3D point cloud feature of an object represents its shape, created by densely gathered points in 3D space. We adopt point clouds to approximate object features as prior information. Our approach involves feature extraction via a point clouds processing network utilizing PointClipV2 [22]. This algorithm facilitates zero-shot classification segmentation to identify features. To reduce computational complexity, the extracted feature map degrades to two dimensions using Principal Component Analysis (PCA). Our visual analysis has substantiated that these 2-dimensional features adequately capture the distinct attributes of various objects, as demonstrated in Fig.6. The distribution of various objects exhibits three distinct directions, leading to their classification into three categories. Elongated objects, like bottles, fall into the first category. Flat objects, such as keyboards, form the second category. Regular objects, like cups, which do not fit into the previous categories, are classified under the third category.
Furthermore, to enhance the algorithm’s ability to generalize across various physical properties of objects, we broaden the training dataset by randomly generating mass and inertia values within a reasonable range for the same object. This approach increases the diversity of physical properties encountered during training. In addition, we reset the environment when the object falls under a threshold or gets out of some bounds. The process enhances the diversity of effective data in the early stage of training.
During the training, we leverage the IsaacGym simulator [31], which enables parallel training, allowing the system to learn the skill of catching various objects simultaneously.
Table 2: Reward Desgin
@@
  • 1
    The reward is composed of five parts, involving the position, posture of the object, task success or failure, and smooth actions.
Appendix C Reward design
@@
The reward is composed of five elements: 1) $r_{p}$ quantifies the distance between the position of the goal object $p_{g}$ and the current object $p_{c}$. 2) $r_{o}$ represents the angular discrepancy between the target object’s orientation and the current object’s orientation, utilizing quaternions for accurate angular measurement. 3) $r_{a}$ evaluates the smoothness of joint movement through the current actions $a_{c}$. 4) A constant reward $c_{1}$ is given for successfully completing tasks $r_{succ}$. 5) A constant penalty $c_{2}$ is imposed for task failure $r_{fall}$. Specifically, the smoothness of joint movement is controlled by negatively rewarding the sum of squares of action quantities as per the strategy. The overall reward is expressed $r_{total}$ in Table 2, where target position distance is calculated using Euler distance, angular distance is determined using quaternions, joint smoothness is penalized using action quantities’ sum of squares, and success and failure rewards are assigned as positive and negative constants, as outlined in Table 2.
Appendix D Details of Lyapunov Stability
We design a stricter constraint: $L_{f_{\pi_{\theta}}}V_{L}(s_{t})<-kV^{\alpha}_{L}(s_{t})$, where $\alpha$ and $k$ are constants ranging from 0 to 1. Note that $kV^{\alpha}_{L}(s_{t})$ is non-negative. This constraint avoids the issue of the Lyapunov function having very small values, thus speeding up the convergence of the stability part in policy improvement.
Our idea is inspired by a classical theorem on finite-time stability convergence in a continuous-time system [32].
Lemma D.1 (Finite-Time Stability Convergence).
In a continuous-time system, the system can be stable within a finite time $T\leq\frac{V_{L}^{1-\alpha}(s_{0})}{k(1-\alpha)}$, if the following condition holds.
$\dot{V}_{L}\leq-kV_{L}^{\alpha}$ (9)
Note that $s_{0}$ is the initial state, $\dot{V}_{L}$ is the Lie derivative $L_{f}V_{L}$, $k$ and $\alpha$ are constants ranging from 0 to 1.
Proof.
First, we perform an integral operation on both sides of the condition.
@@ Appendix D Details of Lyapunov Stability
Then, we can obtain
$\begin{split}\frac{1}{\alpha-1}\frac{-1}{V_{L}^{\alpha-1}}|^{T}_{0}&\leq-kt|^{T}_{0}\\ \frac{1}{\alpha-1}(V_{L}^{1-\alpha}(s_{0}))&\leq-kT\\ \frac{V_{L}^{1-\alpha}(s_{0})}{1-\alpha}&\geq kT\end{split}$ (11)
where $s_{0}$ is the initial state. Finally, the convergence time $T$ can be represented as:
$T\leq\frac{V_{L}^{1-\alpha}(s_{0})}{k(1-\alpha)}$ (12)
Compared with $\dot{V}_{L}\leq 0$, the following lemma accelerates the convergence time from infinity to finite time. We can provide an intuitive explanation. Equation 9 guarantees that the derivative of the Lyapunov function $V_{L}$ is constantly less than a large negative value ($-kV_{L}^{\alpha}$) instead of 0. Thus, we can similarly use the condition of Equation 9 in our method. This process not only facilitates the policy in finding behaviors with better stability properties but also avoids the values of the Lyapunov function being close to 0 in the entire state space. For the choice of $\alpha$ and $k$, we set $\alpha$ to 0.7 and increase $k$ linearly from 0.1 to 1 according to the training steps. The small $k$ in the early stage can alleviate the collapse of the model.
Appendix E Ablation Experiment Analysis
The inclusion of Lyapunov stability ($a.$) in the baseline PPO notably enhances the hand’s catching ability, particularly accentuated during complex tasks like the Overarm Catch. By guiding dexterous hands to avoid unstable gestures and focusing on learning stable catching motions, we achieve higher success rates and rewards. It is worth noting that in the simplest task, the Underarm Catch, the effect of Lyapunov stability is limited. As seen in Fig.3, the task’s stability region is large, thereby making it hard to balance between stability and rewards. Compared with PPO, it is more challenging to explore optimal solutions for success rate.
The impact of increasing intrinsic advantage ($b.$) is observed primarily during the early stages of learning. While its effect may not be substantial, it demonstrates the capability to expedite early learning and convergence across multiple tasks. Parameter normalization ($c.$) significantly enhanced the ability of the throwing process. The normalization layer accelerates the early convergence of the model due to a more stable gradient. The stable gradient update benefits searching the optimal point, particularly in tasks with relatively simple throwing processes. It performs well in tasks where the success of the complete throwing-catching task depends heavily on the throwing process, such as Under2Overarm Catch, Underarm Catch, and Abreast Catch. The incorporation of the object’s point cloud feature ($d.$) simultaneously considers both the throwing and catching processes, offering valuable cues for each. This enhancement notably boosts the algorithm’s performance by improving convergence speed and task success rates. Given that the object’s shape affects the catching, this feature significantly improves tasks with challenging catching behavior like Overarm Catch. Because the throwing-catching behavior is insensitive to the object’s shape in Underarm Catch, adding point cloud to extend state space harms the performance slightly.
Appendix F Generalization Ability
For diverse objects: In this section, we delve into a detailed analysis of our method’s generalization capabilities. It demonstrates the ability to execute distinct catching gestures based on the positions of the shadow hands and the objects being caught. Through an analysis of the high-dimensional point cloud features, we categorize most objects into three primary classes. Remarkably, our method exhibits over three different catching gestures for each of these three object types, as depicted in Fig. 7. For objects falling into the regular, flat, and elongated categories, the model generalizes three distinct gestures in each of the five tasks. For instance, when dealing with regular objects like apples, it employs either a half-grip or pinching gesture. Slender objects like bananas are handled with a full-grip position. Flat objects, such as scissors and pokers, are directly held and secured between the fingers and palms. This versatility in catching gestures reflects the algorithm’s remarkable adaptability to different object types and positions.
Appendix G Simulation to Real World Design
To ensure the successful transfer from the simulation environment to the real world, we address key challenges associated with sim-to-real transfer and offer specific solutions to overcome them.
Challenge 1: Can the observations obtained during the simulation process be effectively acquired in the real world?
Solution 1: In the following, we introduce how to obtain each component of the observation in the real world.
  • Position, rotation, velocity, target position, rotation of the object. We have validated the use of RGB image input from the main viewing angle, utilizing GPT-4V or other multimodal large models in combination with object detection and segmentation technologies. This approach enables us to obtain the initial pixel position of both the current object and the target. Real-time 6D pose estimation allows us to track the object’s pose and pixel position during its flight. By integrating data from a binocular or depth camera, we can extract the object’s three-dimensional world coordinates, and by leveraging the video stream frame rate, we can calculate the object’s motion speed.
  • Point cloud prior information of the object. The object point cloud feature serves as prior information, capturing the shape characteristics after standardizing the object’s size. This feature does not need to be acquired in real-time during the process. When acquisition is not feasible, the point cloud feature can be substituted with the representative geometry of the three types of objects detailed in Appendix B.
  • The position, rotation, and speed of shadow hands’ wrists and joints. These can be obtained from the robot itself.
  • The force of the fingertips. These can be obtained through a fingertip force sensor.
Challenge 2: How can we address the observation errors that occur in the real-world observations?
Solution 2: To account for potential errors in real-world observation acquisition, we introduce a pose perturbation of ±5° on each of the three axes of the current hand posture and a perturbation of ±5 cm in the distance between the hands. This approach is designed to test the policy’s tolerance for variations in hand positions and postures. Additionally, we incorporate an observation error of ±5 cm and a pose perturbation of ±20° on each axis for the position and pose measurements of the object and target, respectively, along with an observation error of ±20% for the object’s velocity measurement. The success rate measured after applying all these errors is compared with the original success rate, as presented in Table 3.
Figure 9: Video snapshots of throwing-catching tasks performed by our method. The description of tasks can be found in Appendix G.1.
Table 3: Success rate of Task with disturbance
@@ Appendix G Simulation to Real World Design
Orignal(11 object) Add disturbance (6 object)
META(Train) META(Test) META
Overarm 60.30 69.97 60.17
Overarm2Abreast 68.77 79.88 66.75
Under2Overarm 73.95 71.06 70.09
Abreast 72.95 79.97 68.84
Underarm 88.99 73.51 83.65
Challenge 3: Can the action outputs from the simulation process be implemented for control in the real world?
Solution 3: To ensure that the actions generated in the simulation can be effectively controlled in the real world, our algorithm framework utilizes most position-based control, and minimizes reliance on force-based control. In our approach, the finger joints of the shadow hands are position-controlled, while only the force and torque of the wrists are controlled. When the robotic arm is connected to the wrist and transitioned to the real-world setting, the planning and control of the robotic arm can be handled by an impedance controller after getting the force and torque of the wrists. This design facilitates future research and verification tasks involving the connection of the wrists to the robotic arm.
Challenge 4: Can our method generalize to different hand poses with significant variations?
Solution 4: Although we believe that the five hand postures designed in this work already include the main combinations of the three axes of the two-hand xyz, and the other hand postures can be varied from these hand postures, in order to ensure the effectiveness of our policy, we added three more general postures: overarmout45, abreastin45, and overarmrightleft45, and achieved a good meta-success rate on the test set and the training set after training, as shown in Table 4. A schematic diagram of the task is shown in Figure 9.
Table 4: Success rate of New Task
[…]
META
Overarmout45 61.44
Abreastin45 68.77
Overarmrightleft45 85.41
Finally, regarding future work, we will support further research on the combination of dexterous hands with robotic arms.
G.1 Description of Added Tasks
The new tasks are depicted as follows.
  • Overarmout45 Catch: The hands are oriented vertically. Then, the throwing hand rotates around the x-axis of the reference coordinate system (shown on the right side of Fig. 9) by -45 degrees, while the catching hand rotates around the x-axis by 45 degrees.
  • Abreastin45 Catch: The setting of hands is built upon the Abreast Catch task. Then, the throwing hand rotates around the z-axis of the reference coordinate system (shown on the right side of Fig. 9) by 45 degrees, while the catching hand rotates around the z-axis by -45 degrees.
  • Overarmrightleft45 Catch: The hands are oriented vertically. Hereafter, the throwing hand rotates around the y-axis of the reference coordinate system (shown on the right side of Fig. 9) by 45 degrees, while the catching hand rotates around the y-axis by -45 degrees.
Appendix H Comparison Experiments Additional Notes
To further demonstrate the effectiveness of the proposed method compared to baseline and other reinforcement learning approaches, we conducted an additional comparative experiment, focusing on the success rate, as shown in Table 5. The results clearly indicate that our approach consistently outperforms both the baseline and other methods across all metrics.
Table 5: Success rate of Task with Other Algorithm
[…]
LTC PPO TRPO SAC DDPG
META(Train) META(Test) META(Train) META(Test) META(Train) META(Test) META(Train) META(Test) META(Train) META(Test)
Overarm 60.30 69.97 5.16 0.71 0.84 0.96 0.00 0.39 0.58 0.65
Overarm2Abreast 68.77 79.88 24.75 34.41 0.00 0.54 0.00 0.00 6.81 5.51
Under2Overarm 73.95 71.06 62.76 66.34 14.88 19.12 68.82 69.86 53.85 59.24
Abreast 72.95 79.97 43.08 52.76 47.91 44.88 0.00 0.00 0.00 0.00
Underarm 88.99 73.51 70.74 65.81 4.01 4.425 0.00 0.00 0.00 0.00